snowflake0531:

What is the amplitude, period, and phase shift of f(x) = −4 sin(2x + π) − 5?

snowflake0531:

darkknight:

amplitude is the number in front, aka 4 the period is how much time it takes for one revolution, u can graph this if u wanna, the phase shift is i think how much up and down? that would be how much you shift the function up and down

darkknight:

@AZ ? i cant really think rn

snowflake0531:

well i just know it's either the first or second too

darkknight:

forgot the negative, my bad for the amplitude but pretty sure amplitude always positive

darkknight:

yea im being dumb phase shift isnt up or down, wait for AZ

snowflake0531:

lol

AZ:

y = A sin(B(x + C)) + D A is our amplitude \(\text{period} = \dfrac{2\pi}{B}\) phase shift is C. If it's +C then it's going to the left. If it's -C then it's going to the right D is our vertical shift

snowflake0531:

So which one is it, like, how is the period and phase shift determined then

snowflake0531:

okay then soooo, explain how to do it please?

AZ:

y = A sin(B(x + C)) + D is what I had written The equation they gave you is like this \( y = A \sin(Bx + BC) + D\) \(f(x) = -4(\sin(2x + \pi)) - 5\) So first, let's factor out the 2 from \(2x + \pi\) so we can find the phase shift and period

darkknight:

hmm but amplitude is how far from the center,  sorry if im being dumb xD

snowflake0531:

so.... AZ how do i figure out phase shift and period

snowflake0531:

can i guess the second one

AZ:

I think you're right, darkknight. Amplitude should be the positive value regardless :)

snowflake0531:

okay, i kinda need to turn this in soon, could you please explainnnnnnn what the answer is pleaseee

XioGonz:

f(x) = A sin (B(x + C)) + D f(x) = -4 sin (2x + π) - 5 then take 2 from (2x + π) f(x) = -4 sin 2(x + π/2) - 5 A = 4 , B = 2 C = π/2 D = -5 A = the amplitude Amplitude = -4 Period = 2π/B The period = 2π/2 = π C is the phase shift The phase shift π/2 to the left Amplitude = -4 period = π phase shift: x = π/2 SO Amplitude = -4; period = π; phase shift: x = π/2 is your answer

XioGonz:

LOL I had forgotten to post

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 so.... AZ how do i figure out phase shift and period \(\color{#0cbb34}{\text{End of Quote}}\) So we have \(2x + \pi\) Let's factor out 2, what do we get? \( 2(x + \dfrac{\pi}{2})\)

AZ:

Now your original equation was \( f(x) = -4(\sin(2x + \pi)) - 5\) We factored out the 2 so we have \( f(x) = -4(\sin(2(x + \dfrac{\pi}{2})) - 5\) So now it's in the same form as this \(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ y = A sin(B(x + C)) + D A is our amplitude \(\text{period} = \dfrac{2\pi}{B}\) phase shift is C. If it's +C then it's going to the left. If it's -C then it's going to the right D is our vertical shift \(\color{#0cbb34}{\text{End of Quote}}\)

AZ:

What is the phase shift of \( f(x) = -4(\sin(2(x + \dfrac{\pi}{2})) - 5\)

snowflake0531:

thank you xiogonz and i guess it would be x/2, AZ

XioGonz:

Np

AZ:

It's \(\Large\pi\) not x haha

snowflake0531:

wait i got it wrong

darkknight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ What is the phase shift of \( f(x) = -4(\sin(2(x + \dfrac{\pi}{2})) - 5\) \(\color{#0cbb34}{\text{End of Quote}}\) when you do something like sin(x+1) that ^^^ actually shifts 1 to the left

darkknight:

in that context, if you shift +pi/2 it wouldnt shift to the right, so that is not the phase shift

AZ:

The font size, even in your question, is pretty small so the period looks like it's listed as x but it's \(\pi\)

snowflake0531:

ye lol

AZ:

looks like an n* but it's actually \(\pi\)

darkknight:

oh breh, amplitude always positive, also \(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight \(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ What is the phase shift of \( f(x) = -4(\sin(2(x + \dfrac{\pi}{2})) - 5\) \(\color{#0cbb34}{\text{End of Quote}}\) when you do something like sin(x+1) that ^^^ actually shifts 1 to the left \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight in that context, if you shift +pi/2 it wouldnt shift to the right, so that is not the phase shift \(\color{#0cbb34}{\text{End of Quote}}\)

snowflake0531:

---welp

AZ:

Yes, since it's shifting to the left it's going to be negative

AZ:

and amplitude is always positive

darkknight:

@snowflake0531

snowflake0531:

i think i forgot how to do these things again o-O

snowflake0531:

at least i still have my notes 😔